Dr. Mack Lemore, an expert in consumer behavior, wants to estimate the average amount of money that people spend in thrift shops. He takes a small sample of 8 individuals and asks them to report how much money they had in their pockets the last time they went shopping at a thrift store. Here is the data: 17.24, 20.63, 26.76, 19.85, 18.24, 23.38, 29.55, 28.29. He wishes to test the null hypothesis that the average amount of money people have in their pockets is equal to $20. Calculate the test statistic to two decimal places. Take all calculations toward the answer to three decimal places.
Answer #1
Values ( X ) Σ ( Xi- X̅ )2
17.24 33.0913
20.63 5.5814
26.76 14.1941
19.85 9.8753
18.24 22.5863
23.38 0.1502
29.55 43.0008
28.29 28.0635
Total 183.94 156.5429

Mean X̅ = Σ Xi / n
X̅ = 183.94 / 8 = 22.992
Sample Standard deviation SX = √ ( (Xi - X̅ )2 / n - 1 )
SX = √ ( 156.5429 / 8 -1 ) = 4.729

Test Statistic :-
t = ( X̅ - µ ) / ( S / √(n))
t = ( 22.992 - 20 ) / ( 4.729 / √(8) )
t = 1.79